OpenStudy (anonymous):

Sandra is riding the Ferris wheel, and her height can be modeled by the equation H(t) = 25 cos pi/14 t + 31, where H represents the height of the person above the ground in feet at t seconds. Part 1: How far above the ground is Sandra before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Sandra begins the ride at the top, how far from the ground is the edge of the Ferris wheel when Sandra's height above the ground reaches a minimum?

OpenStudy (anonymous):

Is it \[H(t)=25\cos(\frac{ \pi }{ 14 }+31) or H(t)=25\cos(\frac{ \pi }{ 14 })+31\]

OpenStudy (anonymous):

\[H(t)=25 \cos \frac{ \pi }{ 14 } t+31\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Part A is just the y-intercept.

OpenStudy (anonymous):

whats the y intercept then? i dont know what the number is

OpenStudy (solomonzelman):

set t=0, and solve for H(t)

OpenStudy (anonymous):

so its \[H(0) 25\cos \frac{ \pi }{ 14 } (0)+31?\]

OpenStudy (anonymous):

what about part 2? @SolomonZelman

OpenStudy (solomonzelman):

The circumference of a wheel is a revolution.

OpenStudy (anonymous):

so whats the equation to figure out the circumference and to find out how long it takes to make a rotation

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/hd9hwovwja And the equation is \(\LARGE\color{blue}{ H(t)=25\cos(\frac{π}{14}t)+31 }\) Looking at the graph it is 28 units on the x-axis, to get the full curve looking like • • • • • • • • from 0 to 28 • • • • • • • •

OpenStudy (solomonzelman):

I am not sure on this one

OpenStudy (solomonzelman):

And for part 3, it is asking for the distance from the maximum to the minimum point ( I think )

OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: f(x) g(x) f(x) = 2 tan(3x + π) g(x) = 3 sin(4x - π) - 2 Using complete sentences, explain how to find the y-intercept for each function and determine which function has the largest y-intercept.

OpenStudy (solomonzelman):

set x equal zero and solve for the range (in each function)

OpenStudy (anonymous):

okay? then what?

OpenStudy (solomonzelman):

Solve for the intercepts

OpenStudy (anonymous):

f=.10977 g=-1.83558?

OpenStudy (solomonzelman):

f(x) = 2 tan(3x + π) g(x) = 3 sin(4x - π) - 2 f(0) = 2 tan(3(0) + π) g(0) = 3 sin(4(0) - π) - 2 f(0) = 2 tan( π) g(0) = 3 sin(- π) - 2 f(0) = 2 (0) g(0) = 3 (0) - 2 f(0) = 0 g(0) = - 2

OpenStudy (solomonzelman):

Which function has a larger intercept ?

OpenStudy (anonymous):

f(0)?

OpenStudy (solomonzelman):

YEah